Q
0.100 mol SO2
and 0.100 mol O2
are introduced into an
evacuated 1.52 L flask at 900 K. When equilibrium is
reached, the amount of SO3
found is 0.0916 mol. Use this
data to determine Kp for the reaction.
2SO2
(g) + O2
(g) 2SO3
1
Expert's answer
2019-11-18T06:29:05-0500
"2SO_2 + O_2 = 2SO_3"

"[SO_3] = 0.0916mol\/1.52L = 0.0603 M"

"[SO_2] = 0.100mol\/1.52L - 0.0603M = 0.0055M"

"[O_2] = 0.100mol\/1.52L - 0.0603\/2M = 0.0356M"


"K_c = {[SO_3]^2 \over [SO_2]^2[O_2]}"

"K_c = {(0.0603)^2 \over (0.0055)^2(0.0356)} = 3.367*10^3"

"K_p = K_c(RT)^{2-2-1} = K_c(RT)^{-1} = {K_c \over RT}"

"K_p = {3.367*10^3 \over 8.314*900} = 0.4500"

Answer: 0.4500